PRACTICE QUESTION 1 PT.1: From Coca-Cola’s 10k you have the following information. Prepare a common size balance sheet. SHOW YOUR WORK!

Balance Sheet                             2019               2018           2019                      2018

(in millions)                                                                  Common               Common

                                                                                           Size % ? Size %?

Long term debt                       27,516            28,782     _______                _______                                                                                  

Other liabilities                       37,787            35,376     _______                _______                                                                                  

Total Liabilities                      65,283            64,158     _______                _______                                                                                  

Total Equity                            21,098            19,058     _______                _______

PRACTICE QUESTION 1 PT. 2: Prom the following information, extracted the following from Coca Cola’s 10-K, prepare an indexed comparison of the years. SHOW YOUR WORK!

(In millions)

Year Ended December 31,       2019          2018       2017            2016           2015

Net operating revenues            37,266       34,300    36,212         41,863        44,294

Index number                            _____        _____     _____          _____         _____
Answer #1

Answer

1

Common size
Particulars 2019 2018 2019 2018
Long term debt $      27,516 $      28,782 31.9% 34.6%
Other liabilities $      37,787 $      35,376 43.7% 42.5%
Total liabilities $      65,283 $      64,158 75.6% 77.1%
Total equity $      21,098 $      19,058 24.4% 22.9%
Total liabilities and equity $      86,381 $      83,216 100.0% 100.0%

Long term debt

2019=27,516/86,831=31.95

2018=28,782/83,216=34.6%

Other liabilities

2019=37,787/86,831

2018=35376/83216

Total liabilities

2019=65283/86381

2018=64158/83216

Total equity

2019=21,098/86381

2018=19,058/83,216

Same formula is used for others.

Q R S T U 2019 27516 Particulars Long term debt Other liabilities Total liabilities Total equity Total liabilities and equity

2

Index number 100 is assumed to be appointed to 2015 Revenue

Index number for 2016=Base index number*2016 Revenue/2015 Revenue

=100*41,863/44,294

Index number for 2016=94.51

Index number for 2017=Base index number*Revenue 2017/Revenue 2015

=100*36,212/44,294

Index number for 2017=81.75

Index number for 2018=Base index number*Revenue 2018/Revenue 2015

=100*34,300/44,294

Index number for 2018=77.44

Index number for 2019=Base index number*Revenue 2019/Revenue 2015

=100*37,266/44,294

Index number for 2019=84.13